OpenStudy (anonymous):

How do i plot the graphs of : 1. y=Asin(kx+wt+pie/6) 2.y=Asin(kx+wt-pie/6) 3.y=Asin(kx-wt+pie/6) 4.y=Asin(kx-wt-pie/6)

OpenStudy (anonymous):

somebodyy pls help!

OpenStudy (anonymous):

well you can't cause the equation takes in 2 inputs x and t, so there is no way you can draw a graph which depicts how the y is changing as both x and t are changing.. i know that is a wave equation, usually we draw the graph of this at SOME time interval for example.. consider first one.. and let say t=0. How would the wave look like? well the quation becomes y= Asin(kx+pi/6) when x = 0 y= Asin(pi/6)= A/2. hence it would look something like its difficult to draw with mouse.. i shall use my pen tablet and redraw it!